UT 2021 Oct 27 - D/E Night



Night Info:

  • Split night between LBTI AO-onsky checkout, D/E tasks.
  • Night Center: 00:04
  • Handover from LBTI to LUCI-2 checkout: 23:59
  • Moon rise: 21:48
  • Sunset: 17:39
  • Twilight ends: 18:58
  • Twilight starts: 05:10
  • Sunrise 06:29


Night Plan:

  • LBTI on-sky checks until 23:59
    • thin clouds
    • seeing < 1.2"
  • LUCI-2 checkout starts 00:09
    • DONE Rotator ZP - via astrometric masks - 0.5hr
      • thin clouds, seeing < 1.2"
      • Stone O (from UT 2021-10-07) - theta = -0.01 star#36 rejected)
      • Stone P sets 00:15 (30 deg)
      • Stone A sets 02:20 (30 deg) - worked best without any stars rejected for LUCI-1
      • Could test other LUCI MOS masks for extra datapoints or LS test
      • Files needed: ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1
        • /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
        • Bounce GCS and PCS
    • AGW Transform - 1-1.5hrs
      • thin or variable clouds, seeing < 1.2"
      • Transform Collect
      • Potential lists:
        • Stone_A3_LUCI.list - sets 02:20
        • Stone_C1.list - RA ~04hr Dec -01 - sets twilight
        • Stone_D_new_Oct27.list - RA ~04:30hr Dec -00 - thru twilight
        • Stone E (-90,-135,90, 0 PAs) - RA 07hr Dec +00 - 01:30hr-twilight
    • Photometric standards bino-LUCI if poor seeing but clear
  • MODS
    • DONE IT 8472 - GS at left edge of patrol field - 0.25hr
      • any conditions, but see below if photometric
      • Run sdss04+51_phot.txt in /home/lbto/Service_Observing/2021B_D_E/MODS
      • Takes 30-sec exposures at g+u(2x) and riz.(need star bright enough to WFS/guide on). We could make use of photometry since the field is in SDSS DR12.
  • LBCs
    • DONE Active TMS - 3hr
      • Instructions
      • any conditions
      • start the night with a previous reference (the reference from 20211014), note the corrections indicated by dohybrid.
        • reference is tms?x_ref.dat_ and is in telescope@robs.mountain.lbto.org:~/TMS_logs/20211014/
        • on robs as user telescope, cp ~/TMS_logs/20211014/tms_?x_ref.dat /tmp
      • track a field through a large elevation change for ~2 hrs
        • Choose GD_279 or J031900+43. The scripts will dither 6x and change filters. You would need to repeat the OB about 4-5 times to make up 2 hours on the field.
        • visibility plot: 2021_oct26_2stds_vis.png
        • Scripts in: /home/lbto/Service_Observing/2021B_D_E/LBC/New_PhotFields/Scripts
        • Dual purpose: Please note whether it is photometric or not since these fields are from the set of new phot stds.
      • slew to a few fields at different elevations with TMS in active mode. Run FPIA and note the corrections indicated.
        • Use copointing fields for new phot standards.
        • Want to change elevation (see vis plot png or OT New_Phot_Fields_Master_OT_V1.1.xml).
          • visibility plot, 2021_oct_26_stds_vis.png.
  • Science Programs:
  • Other Poor Weather:

NightLog:

UT 00:02 - starting LUCI GUIs -init all & test darks
LUCI-2 MOS Error

UT 00:06 - cannot initialize pre-condition violation - no mask in use at start

-ok strain gauges are on - as seen by the Switchbox

-zeroing the strain gauges

-initializing the MOS

L1 0001-0002 - 2.51sec dark

L2 0021-0022 - 2.51sec dark

putting LUCI-1 mask away. Not likely to need a mask tonight for LUCI-1

UT 00:16 - Bringing up LBC GUIS

UT 00:39 - prepping to open with LBTI...

LBTI

UT 00:44 - enclosure open - LBTI starting....

UT 01:04 - bringing up MODS on OBS2 - MODS stuff isn't readily available on OBS5 - im not updating the paths if thats the issue - it should b ready

UT 01:07 - handing over to facility instruments....

MODS Standard Stars

UT 01:09 - taking simsnap

UT 01:13 - reconfiguring for MODS

-start with M71

UT 01:28 - sending pointing preset

UT 01:39 - seeing is only 2.3" on the guiders
M71

UT 01:41 - sending prest to M71 Dual

seeing on guiders 2"-2.2"
  • Airmass = 1.052
  • M1B 0003-0004
  • M1R 0003-0006
  • M2B 0003-0004
  • M2R 0003-0006
UT 01:51 - sending preset to M71 Red
  • Airmass = 1.058
  • M1R 0007-0009
  • M2R 0007-0009
UT 01:58 - sending preset to M71 Blue
  • Airmass = 1.066
  • M1B 0005-0006
  • M2B 0005-0006
Guiders bouncing from 1.8"-2.4"
SDSSY3

UT 02:04 - sending preset to SDSSY3 Dual
  • Airmass = 1.307
  • M1B 0007-0010
  • M1R 0009-0014
  • M2B 0007-0010
  • M2R 0009-0014
Guiders showing 1.9"-2.2", hit 2.6" on MODS2

UT 02:19 - there was a definite pause in M1B between the first and 2nd u-image, not sure why.

UT 02:21 - long read out of M1R - still showing "Reading Out..." - reading out 1min later

UT 02:23 - sending preset to SDSSY3 Red
  • Airmass = 1.248
  • M1R 0015-0020
  • M2R 0015-0020
UT 02:38 - sending preset to SDSSY3 Blue
  • Airmass = 1.201
  • M1B 0011-0014
  • M2B 0011-0014
UT 02:43- seeing is so bad we can't measure it well in the guider/wfs.....i'm guessing 0.08" FWHM in the WFS is a bit off...
SDSSY_r015

UT 02:49 - sending preset to SDSSY_r015 Dual

UT 02:52 - fuzzballs too big couldn't lock on - resending preset

UT 02:57 - little struggle to acquire but seeing now 2.2"-2.4"
  • Airmass = 1.792
  • M1B 0015-0018
  • M1R 0021-0026
  • M2B 0015-0018
  • M2R 0021-0026
UT 03:06 - sending preset to SDSSY_r015 Red
  • Airmass = 1.694
  • M1R 0027-0032
  • M2R 0027-0032
UT 03:08 - AGW5 is a champ, grabbed that guidestar on the edge and pulled it right in 2x

UT 03:10 - Resending MODS2 script - timed out first try

UT 03:11 seeing is 2.3"-2.4" on the guiders

UT 03:18 - MODS1R slow readout again - #0032

UT 03:21 - sending preset to SDSSY3_r015 Blue - MODS1 (M2 is out of sync due to earlier abort)

-failed to find star.

UT 03:22 - pointing check

UT 03:26 - sending preset to SDSSY3_r015 Blue

MODS1 Error:

STATUS: WARNING: **FAULT** - BGRATING BGRATING=TIMEOUT cannot read from 192.16DONE: INSTCONFIG completed with 1 faults

Using troubleshooting steps: https://wiki.lbto.org/Instrumentation/TroubleshootingISS

Runing pingMODS1 - everything came back online

Reset microLynx controller #9 (this is blue grating)

isisCmd --mods1 m1.ie ieb b mlc 9 off

isisCmd --mods1 m1.ie ieb b mlc 9 on

Then reset it:

lbto@obs2:50 % isisCmd --mods1 M1.ie bgrating reset
*** ERROR: BGRATING BGRATING=FAULT RESET Failed, seek did not reach valid position, in-position sensor did not assert
on OBS-2:

lbto@obs2:52 % isisCmd --mods1 m1.ie bgrating
DONE: BGRATING BGRATING=1 GRATNAME='Flat'
Clicked Update on GUI and it went away....(error)

Clicked Grating and COMMIT on M1 GUI...

and back to imaging OK

UT 03:38 - resending blue preset to SDSSY3_r015
  • Airmass = 1.471
  • M1B 0019-022
  • M2B 0019-022

-also can run through Feige 110 at multiple airmasses, and or BD+28 spectrostds

MODS SpecPhot Std

BD+284211

UT 03:49 - sending preset

Seeing 2.4-2.5"

MODS1 Offset Command:
offsetxy -1.104 12.060 rel
MODS2 Offset Command:
offsetxy 4.078 8.153 rel

Tweaks:

MODS2 Offset Command:
offsetxy -0.453 0.455 rel
MODS1 Offset Command:
offsetxy -0.492 0.248 rel

UT 03:59 -starting spectroscopy - running through all modes
  • Airmass = 1.039
  • M1B 0023-0034
  • M1R 0033-0046
  • M2B 0023-0034
  • M2R 0033-0046
UT 04:16 - seeing 1.7-2.2"

-will run through prism mode too

UT 04:38 - seeing down to 1.2-1.7"
Feige 110

UT 04:48 - sending preset

UT 04:50 - seeing 1.7"-1.8"

MODS2 Offset Command:
offsetxy 3.505 8.573 rel
MODS1 Offset Command:
offsetxy -1.355 11.590 rel
Tweaks:

MODS2 Offset Command:
offsetxy -0.577 -0.108 rel
MODS-1 looks good no tweaks.
  • Airmass = 1.277
  • M1B 0035-0046
  • M1R 0049-0060
  • M2B 0035-0046
  • M2R 0049-0060
UT 04:59 - seeing back up to 1.9"-2"

UT 05:06 -MODS 1R slow on readout again, finally at UT 05:09

MODS2Blue reports shutter open when doing Red Grating, but MOD1Blue shows shutter Closed

UT 05:18 - seing down to 1.48-1.6"

UT 05:20 - GUI on M2B shows shutter closed after setting up for blue grating and IMCS lock. Pehaps this was a GUI glitch?

UT 05:39 - seeing trending downards - 1.2"-1.3" on guiders

-spoke too soon?

IT 8472 Test - WFS cutout Issue

UT 05:47 - sending preset, since its clear will run to completion - can be used for standard stars too

-no evidence of a cutout problem

UT 05:52 - seeing in guiders now 1.1"-1.2' I think we may switch to LUCIs to get the RotZP and Transform

LUCI Checkout

UT 05:58 - reconfiguring for LUCIs

-start with Stone P - putting MOS mask in FPU to save time while we reconfigure
LUCI-2 MOS Error

Error moving Mask 18 to FPU El 90 Rotator = 210

Post Condition Violation Exception MHU absolute angle

NOT True that motor is at required position

picker arms open - everything is locked - this was going to grab the mask.

Position 18 -12888 current position -12887

Translator angle should be 24.3, current position is 24.67

Initializing the MOS since nothing is open nothing unlocked --> move the robot to a knonwn position

Cleared it - moving to grab mask again

UT 06:12 - config to LUCIs done
LUCI-2 Rot ZP

UT 06:14 - pointing check
StoneP

-seeing on guider is borderline, but its also low - will give it a try and move to next MOS

#36 looks like worst fit in RA - removing it

#33 next worst, removing it (also faint), removing #35

0.76", -2.28" theta = -0.028 err 0.055, 0.047 Seeing on Guider is 1.2"-1.5" during acquisition

Looks Good
StoneA

UT 06:33 - moving to StoneA

Guider is 1.38", WFS 1.16" - borderline again

seeing down to 1.09" on guider, 1" on WFS

Good fit with all points

0.399", -2.904" theta = -0.050 err = 0.055, 0.055 guider ~1"

Looks Good

NO UPDATE TO ROTZP Needed
AGW Transform

UT 06:48 - putting mask away Acquire preset to Stone A3

UT 06:51 - sending an active preset to tune up collimation before we do this...

-putting N30 fieldstop into place w/N375 camera, H-filter DIT=4sec, flexure off

L2 0031 - test exposure

UT 06:56 - starting collect - seeing may have blown up when I clicked 'enter'....

UT 07:01 - GCS shows FWH=3" after 2nd exposure
  • L2 # GS
  • 0032 0
  • 0033 - off image
  • 0034 10
  • 0035 5
  • 0036 1
UT 07:07 - Stopping this.....too highly variable with excursions to 2"-3"

LBC - TMS Testing

  • ssh telescope@robs
  • cd /lbt/lbto/supportscripts/MetrologySupport
  • The reference file was copied over to /tmp by J. Hill and O.Kuhn
  • in MetrologySupport directory execute tms_loop.pl once we are configured and authorized
  • scripts already exported, finder charts in sub-directory
  • GD279 or J0319 targets for night
  • Go to co-point field, run dofpia
  • co-point with lbcrangebal, then go to field of choice
  • let screen run - will need to restart script 3-4x
  • when moving to a new field co-pointing scripts run 20seconds at least
UT 07:29 - error on swingarm move - J.Williams going into chamber to check

UT 07:48 - TMS lasers on

UT 07:58 - moving to J0319 - sent co-point script, hit STOP, track preset going

UT 08:04 - winds too high in the north

UT 08:06 - shell RIP'd before we could move to the south

UT 06:08 - moving to SDSS_Y4 01:54:04.4 +14:39:54 2000.

-getting errors about missing channel 11

UT 06:13 - CTRL-C 2x to run DOFPIA
  • wait a few cycles before running dofpia
  • setting reference: ./lbc_set_tms_offset.py -s 2
    • this will take a while before returning prompt
  • ./tms_loop.pl
  • sending co-pointing script
copoint_SDSSY4.png

can't copy and paste from X2GO on OBS5!!!

UT 08:25 - sending script to SDSSY4

UT 08:53 - set #2

UT 08:55 - winds gusted over 22m/s - CTRL-C'd TMS, stoped LBC OB

CLOSING DOME (yes I said DOME!)

Closed Dome LBC Checks

  • 23May2021_Linearity.xml
test with Chamber Foot Lights and 20sec V-Bessell and i-sloan

-2k counts red

test with Chamber Foot Lights and Stairwells

-1800-2000 Blue, 3k counts red

Lower Chamber floods - 1sec

-20k blue, 2k red

UT 09:46 - LBC-R closed dome gain - LBCR_gain_lower_floods.xml

UT 10:14 - data collection complete

LUCI-2 Closed Dome Checks

UT 09:40 - putting N30 fieldstopa away, will take detector readouts for D.Thomson

L2 0038 - 2.51 sec lower chamber floods Ks

L2 0039-0048 - 10 x [10 x 2.51s] LIR/INT Ks N3.75 imaging

UT 09:51 - done with first set - putting blinds in

Resuming Observing

UT 10:35 - winds below llimits - will go to a new field
  • starting the loop again tms_loop.pl
  • heading to field to run dofpia
  • running dofpia
  • do copointing
copoint_SDSSY4_2.jpg

- don't seem to be sure if TMS should be stopped before doing DOFPIA - discussion going on whether DOFPIA and TMS might have fought each other - we ran DOFPIA while TMS active was going on

-the copoint images have coma and look bad, particularly blue
  • CTRL-C 2x TMS
  • Running DOFPIA again
  • PRO TIP: DO NOT KEEP TMS RUNNING WHILE DOFPIA IS RUNNING - PAUSE TMS FIRST
  • Running OB for field
  • Upon arrivel - run setpoint
  • Now run TMS active loop
LBC-B 1.33" FWHM, LBC-R 1.52" FWHML

UT 11:22 - taking set #2

UT 11:36 - losing focus on LBC-R - shows doughnuts
  • -stopping early
  • -CTRL-C out of TMS
  • -running DOFPIA
  • sending preset, hit STOP
  • setting reference point
  • setting TMS active
  • sending preset and letting it go
UT 11:43 - starting set #3

-skirting the wind limits again

UT 12:03 - starting set #4

UT 12:07 - out of focus
  • stopping script and TMS
  • hit CTRL C in the wrong window - OT
  • re-running DOFPIA
  • sending preset and stop
  • set point
  • starting TMS
  • sending preset again - UT 12:21
UT 12:25 - Blue showing a lot of coma, Red looks good

UT 12:26 - stopping and running DOFPIA again
  • sent preset, stopped
  • settingp point
  • starting TMS active
  • sending preset again UT 12:34
UT 12:37 - blue still has coma, red looks OK

UT 123607 - LBC-Blue scrambled

Hardpoint error message on SX Primary - could this be affecting and/or causing the blue coma

LBC-B 124438 - double image - jump? trail?

UT 12:53 - images are saturated in red- script finishing now

UT 12:54 - closing dome

UT 12:55 - powering off LBCs, closing out LUCI GUIs

MODS Calibrations

Dual Imaging Flats

M1B 0050-0059

M1R 0064-0078

M2B 0050-0059

M2R 0064-0078
Blue Imaging Flats

M1B 0060-0069

M2B 0060-0069
Red Imaging Flats

M1R 0079-0093

M2R 0079-0093

-need to run fistflush on M2R for 0087 onwards - OK
3K Bias

M1B 0070-0074

M1R 0094-0098

M2B 0070-0074

M2R 0094-0098

UT 14:21 - putting MODS-2 to sleep

UT 14:26 - putting MODS-1 to sleep


-- BarryRothberg - 20 Oct 2021
Topic revision: r6 - 30 Oct 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback